ELECTRONIC PHOTODISSOCIATION SPECTROSCOPY OF GAS-PHASE IrBr_{n}^{k−}

\underline{JESSE C. MARCUM}, J. MATHIAS WEBER, JILA, NIST, and Department of Chemistry and Biochemistry, University of Colorado, Boulder, Colorado 80309.

Iridium complex anions (IrBr_{6}^{2−}, IrBr_{5}^{−} and IrBr_{4}^{−}) were studied using photodissociation spectroscopy in the photon energy range between 1.1 and 5.6 eV. These spectra probe the electronic structure of iridium bromide complexes as a function of the coordination number and metal oxidation state. The observed spectra for isolated IrBr_{6}^{2−} contain similar features to its spectrum in aqueous solution. An analysis of the fragment channels reveals a tendency for formation of IrBr_{5}^{−} fragment ions in the visible while formation of IrBr_{4}^{−} fragment ions is favored in the ultraviolet. Spectra for IrBr_{5}^{−} and IrBr_{4}^{−} parent ions indicate that their preferred fragment channel is the loss of one or more neutral bromine atoms.